i need a answer.
the question is that i need a 7th grade algebra equation with the answer 29.
but the trick is that you can only use the numbers 1,2,3,4 each ONCE.
YOU CAN USE EXPONETS ALSO.
FOR EXAMPLE:
4-2x3+1=7

Answer by stanbon(75887)

i need a answer.
the question is that i need a 7th grade algebra equation with the answer 29.
but the trick is that you can only use the numbers 1,2,3,4 each ONCE.
YOU CAN USE EXPONETS ALSO.
FOR EXAMPLE:
4-2x3+1=7
------------------
Start with x=29
add 1 on each side:
x+1 = 30
Multiply by 2:
2(x+1) = 60
----
Divide by 3
(2/3)(x+1) = 20
Subtract 4 from both sides:
(2/3)(x+1)+4 = 24
-------------------------
If you were to "solve" this equation you would find that x=29.
-------------------------
Cheers,
Stan